# Are AI Detectors Good Enough? A Survey on Quality of Datasets With Machine-Generated Texts

### v1

@@ 1 Introduction
The quality of the outputs of autoregressive Large Language Models (LLM) has grown tremendously in the last five years, making their output almost indistinguishable from human-written texts Chang et al. (2024). This expanded the application fields of such models, as many routine tasks can be entrusted to them nowadays. However, we may observe multiple cases of overuse of these models, when they are utilised for creating texts that are intended to be written and fact-checked by humans. Misuse could be revealed in the generation of fake news Zellers et al. (2019); Zhou et al. (2023), which can mislead readers of such content. Another example is concern in academic society, where a lot of students complete assignments with LLMs Koike et al. (2024); Ma et al. (2023), making professors grade AI-generated content and undervaluing the purpose of the educational process. The spread of machine-generated fragments and plagiarism in scholarly articles increases rapidly with the growth of chatbots and reaches several tens of percent Liang et al. (2024); Gritsay et al. (2023a). Field researches also showed that in the last year alone, more than 60,000 scientific papers contain evidence of the use of machine generation Gray (2024). As generated texts are gaining rapid and widespread popularity, it is crucial to develop systems able to counter the uncontrolled proliferation of artificial data and signal to the reader that the content they read is generated.
@@
Another concern is that the Web is overflowing with machine-generated content, often of poor quality. The fact that such texts contribute bias to publicly available texts on the Internet, through false facts, hallucinations and spelling errors, is worth considering. Given the current agenda of using trillions of tokens from the Internet to teach new language models, Villalobos et al. (2022) revealed that the data will run out by 2028. That means that the training sets for more advanced language models in the future will include a large amount of generated content, as it already has a place in the current picture of available texts from all over the Internet. It was shown in Alemohammad et al. (2023) that such self-consuming will result in substantial degradation of the model’s abilities. Furthermore, the trend is evolving in such a way that human-written texts on almost any topic will be much harder to retrieve. While for texts dated 10 years ago we are confident that the usage of generation was extremely rare, we cannot state the same for modern texts.
@@
Detectors capable to distinguish human-written texts from AI-generated texts, and whose detection quality can be guaranteed, are required by numerous specialists in many fields. We consider that the key to success in building excellent detectors will be high-quality artificial text collections. In this paper we would like to estimate the quality of the available generated data from competitions and research papers. As we are still in a lack of qualitative detectors, even though we see the claimed metrics on various datasets up to 99.9%, meanwhile in the wild a noticeable decline in performance can be seen. Competitions are a rather good way to suggest attractive challenges for researchers, but the results, obtained in these competitions may look confusing: as the models become more and more advanced, seemingly making the detecting task more challenging, participants of competitions still reach almost perfect scores of metric, bringing up the problem about quality of generated data in the provided datasets, see Appendix  C.
Our contributions are following:
@@
    We systemize information about datasets from articles and competitions, dedicated to the detection of AI-generated content task.
@@ 2 Related Work
AI-generated content detection is usually a text classification task, meaning that the input is a text sequence and the output is a discrete, usually binary class prediction. When the task is binary, the common labels are “AI” or “human”, however in some cases one may be interested in multi-class classification between several language models. The last task is usually called authorship attribution. Finally, there is a recent, more complex task, called hybrid writing detection, when text contains fragments of both authors and the task is to determine the borders when authors change.
Among the approaches to tackle the classification problem Jawahar et al. (2020) is to calculate linguistic Fröhling and Zubiaga (2021), stylometric, and statistical features, as well as using classical machine learning methods such as Logistic Regression or Random Forest as classifiers on these features. Some other techniques include the calculation of internal metrics, such as token-wise log probability or rank Gehrmann et al. (2019). They can be calculated for each token and compared for consistency with the prior context. However, these methods are falling behind approaches that use BERT-based models and keep an eye on the context Ippolito et al. (2020); Gritsay et al. (2022) such as RoBERTa Liu et al. (2019) and DeBERTa He et al. (2021a), that are showing state-of-the-art results on this task Uchendu et al. (2021).
@@
Regarding methods for evaluating the quality of the generated data itself, it has become more common to evaluate with the help of LLMs themselves Xu et al. (2023) . The main advantage of such approach is that it does not require any human reference, unlike ROUGE Lin (2004) or MoverScore Zhao et al. (2019). However, the output of model-evaluator needs to be unified and is not always interpretable. Another approach is suggested by Zhu and Bhat (2020), where the text is evaluated based on several linguistic criteria, such as grammar or coherence.
There are many datasets, developed as benchmarks for AI-generated content detection and it would be infeasible to evaluate them all. We will describe the datasets we used in our analysis and experiments in Section 3. There are a number of works dedicated to summarization of information about those datasets, however, only few works focus on the quality of data in them. Participants of competitions often provide a brief overview of the data in their solution descriptions Shamardina et al. (2022b); Rosati (2022); Gritsay et al. (2023b); Boeva et al. (2024); Zhao et al. (2024); Gritsai et al. (2024); Miralles et al. (2024); Valdez-Valenzuela et al. (2024). In Voznyuk and Konovalov (2024) it was shown that in SemEval Task 8 Wang et al. (2024b) there were a number of generations of poor quality, which made the task for detectors much easier. In this work, we analyse some other datasets, whether they also contain similar sort of generation errors, by diminishing the quality of detectors trained on them.
3 Data
@@ 3.1 Datasets From Shared Tasks
Shared tasks to detect artificially generated fragments appeared several years ago. Participants are suggested to solve some tasks connected with detection, either a binary classification, authorship attribution or hybrid human-AI collaboration. In this work, we focused on competitions in which a binary classification task was present. It should be noted that the studies only present data that were made available for training stage, as not all competitions had gold labels provided for test samples after the end of the competition. All shared tasks contain text in English, unless stated otherwise:
RuATD 2022 Shamardina et al. (2022a) contains human- and machine-written fragments in Russian with wide range of themes.
DAGPap 2022 Kashnitsky et al. (2022) contains human- and machine-written scientific excerpts collected by Elsevier.
AuTexTification 2023 Sarvazyan et al. (2023) provides texts for classification in English and Spanish. The organisers considered five different domains to solve with.
IberAuTexTiification 2024 Sarvazyan et al. (2024) is a continuation of the previous competition. The novelty of this shared task is that it is multilingual (six Iberian languages), multi-domain and multi-model.
Voight-Kampff Generative AI Authorship Verification 2024 Ayele et al. (2024) provides two texts, one authored by a human, one by a machine, and participants were suggested to pick the human one. Test data for this task was compiled from the submissions of another competition’ participants and comprised multiple text genres. For shorter term in future mentions, we will refer to this competition as PAN 2024.
SemEval 2024 Task 8 Wang et al. (2024b) focuses on shifts between domains, generators and languages of generated texts. Among available languages are Chinese, Urdu, Bulgarian, Indonesian, Russian, Arabic, German and Italian, however test set consisted only of texts in English, Italian, German and Arabic.
MGT Detection Task 1  [^1] COLING-25 MGT Detection Task 1 is the continuation and improvement of the SemEval Task 8. Organisers aim to refresh training and testing data with generations from novel LLMs and include new languages.
A basic overview of the analysed datasets is presented in Table 1, and a more detailed description of the data source and the topics presented can be seen in Appendix B.
Dataset Year Language Num. of Texts Generated VS Human Average Length Median Length
RuATD 2022 ru 129k 64.5k / 64.5k 236.86 / 221.47 99.0 / 95.0
DAGPap22 2022 en 5.3k 3.6k / 1.6k 799.45 / 1180.07 680.0 / 1126.5
AuTex 2023 en, es 65.9k 33.1k / 32.8k 315.08 / 297.28 386.0 / 351.0
IberAuTex 2024 es, en, ca, gl, eu, pt 98k 52.5k / 45.4k 1036.92 / 1058.36 981.0 / 1018.0
PAN24 2024 en 15.2k 14.1k / 1.1k 2640.50 / 3007.04 2731.0 / 2868.0
SemEval24 Mono 2024 en 34.2k 18k / 16.2k 2465.12 / 2358.05 2570.0 / 2083.5
SemEval24 Multi 2024 en, ar, de, it 42.3k 22.1k / 20.2k 2217.87 / 2256.67 2270.0 / 2032.0
MGT Task 1 Mono 2025 en 610.7k 381.8k / 228.9k 1448.28 / 1541.18 1208.0 / 1080.0
MGT Task 1 Multi 2025 en, zh, it, ar, de, ru, bg, ur, id 674k 416.1k / 257.9k 1422.74 / 1445.33 1195.0 / 1032.0
Table 1: Statistics of the datasets from the shared tasks.
@@ 3.2 Datasets from Papers
The number of collections with generated content has started to increase with an increasing number of available generators. Quite often, researchers together with a new approach for AI content detection publish a parallel dataset on which they have validated their method. In this paper, we picked collections with human- and machine-generated excerpts that are the most common and cited in other researchers’ publications.
GPT2 Output Dataset 222gpt2-output-dataset consists of text outputs generated by GPT-2 models across various prompts.
@@
M4 (Multilingual, Multimodal, Multitask, Massive Dataset) Wang et al. (2023) is a large-scale dataset designed for training models that can handle multiple languages, tasks, and modalities, making it useful for developing versatile AI systems. It contains texts in Arabic, Bulgarian, Indonesian, Russian, Chinese and Urdu, but we focused only on texts in English.
Dataset Year Language Num. of Texts Generated VS Human Average Length Median Length
GPT2 2019 en 1250k 1000k / 250k 2941.28 / 2616.04 3245.0 / 2459.0
HC3 2023 en 85.4k 26.9k / 58.5k 1010.50 / 680.68 1012.0 / 422.0
GhostBuster 2023 en 21k 18k / 3k 3345.07 / 3391.26 3439.5 / 2911.5
MGTBench 2024 en 23.7k 20.7k / 3k 1595.94 / 3391.26 1226.0 / 2911.5
MAGE 2024 en 436k 152.3k / 284.2k 1138.75 / 1281.88 706.0 / 666.0
M4 2024 en 89.5k 44.7k / 44.7k 1587.62 / 3162.40 1454.0 / 1697.0
Table 2: Statistics of the datasets from the research papers.
@@ 4 Approach
We decided to evaluate all datasets on common setups to see how good standard approaches perform on them. We did not have the goal to obtain the highest score, but rather to compare the performance of the same method on different datasets.
4.1 Classifiers
DeBERTa Classifier. A standard approach for binary classification of documents is to fine-tune in supervised manner some classifier, usually BERT-like model. In our case, we used mDeBERTa He et al. (2021b), which is the current state-of-the-art model for machine-generated text detection Macko et al. (2023).
DetectGPT. The DetectGPT framework Mitchell et al. (2023) introduced novel perplexity-based scoring function, which involves perturbing a text passage, after what log-probabilities between original and modified texts are compared. This score is passed as an input to a classification model, with GPT-2 Radford et al. (2019) as the base model and T5-Large Raffel et al. (2019) as perturbations generator. However, as DetectGPT requires intensive computational costs, we utilized Fast-DetectGPT Bao et al. (2024), that substitutes DetectGPT’s perturbation step with a more efficient sampling step. Another advantage is it can be requested via API and approach is applicable without additional fine-tuning stages.
Binoculars Hans et al. (2024) introduce modified score with perplexities and cross-perplexity of two close language models to separate human and machine texts and is proved to be better than approach with classic perplexity. Also, the approach is applicable without additional fine-tuning stages.
4.2 Topological Time Series calculation
It was shown in Tulchinskii et al. (2023) that if we take the inner dimensionality of the manifold on the set of embeddings, we could separate human-written texts from machine-generated ones. However, in that work authors estimated short texts, and therefore we followed the approach in Kushnareva et al. (2024) with sliding window. Topological Time Series calculate intrinsic dimensions (PHD) of the text within sliding window and they can be used as a feature for detector. Authors conclude that this metric can, in fact, help to differentiate the texts of different origin. To be able to compare datasets between each other, we came up with a score, utilizing KL-divergence. Let $h_{d}$, $m_{d}$ be distributions of intrinsic dimensions for two types of texts from the same dataset, of human and machine origin, then our $\text{KL}_{\text{TTS}}$ is following:
@@ 4.3 Perturbations and Shuffling
Based on the results of text modification studies Sadasivan et al. (2024); Mitchell et al. (2023), which show how small perturbations affect machine reading comprehension systems, we decided to consider this way of possibly assessing the quality of a dataset. The key idea here is that AI models are sensitive to such adversarial changes, unlike humans. We considered two modification ideas: Adversarial Token Perturbation and Sentence Shuffling.
Adversarial Token Perturbation. In this approach we divide the text into tokens and randomly replace the token with a synonym from the WordNet Miller (1994) collection with a probability of 70%. We apply such a technique to each represented class. Using an encoder model, we obtain embeddings for each of the texts in the current dataset. Finally, we measure the average embedding shifts for the classes of human and generated texts. We obtain the embedding shifts using the cosine distance between the embeddings of the original texts and the modified ones. As a result, after modifications we obtain $\Delta_{\text{shift}}$ — the log difference of the average embedding shifts.
@@
Figure 1: Comparisons of embedding shifts after two types of modifications for the HC3 dataset.
4.4 Attention Maps
We hypothesised that some patterns in the attention map might be useful in evaluating the quality of the texts, as in the case of perfect quality of generation, the performance of the model on human and machine texts should be identical. As it is infeasible to evaluate attention maps visually for large amounts of texts, we came up with a statistic that may be useful in grasping some patterns. It was intended to grasp “attention columns” or massive activations Sun et al. (2024) that appear on the attention map. For some texts there is no such pattern observed, for others there’s only one distinct column, then there are cases when two or three distinct columns are seen. Our hypothesis was that attention maps for machine-generated texts should show more columns than for human ones, because there may be some tokens that model does not expect to see.
$Attention_{j}(t)=\frac{1}{|t_{j:|t|}|}\sum_{i=j}^{|t|}Attention(q_{i},k_{j})$
Therefore, until the end of the text, it still attends to such tokens, and we calculated the mean difference between the top-1 and the top-2 values across the dataset. Depending on the number of “attention columns”, this difference will vary, but in general we assume that it should be comparable for texts of both origin.

### v2

@@ 1 Introduction
The quality of large language models (LLMs) has grown tremendously in the last five years, making their output almost indistinguishable from human-written texts  Chang et al. (2024). This expanded the application fields of these models, as many routine tasks can be entrusted to them nowadays. However, they can be used for creating texts that are intended to be written and fact-checked by humans. Example of such misuse is generation of fake news Zellers et al. (2019); Zhou et al. (2023), which can mislead readers of generated content. Teachers raise another concern, as many students complete assignments with LLMs Koike et al. (2024); Ma et al. (2023), undervaluing the purpose of the educational process. Machine-generated fragments also appear in academic articles more often with the growth of chatbots and reach several tens of percent  Liang et al. (2024); Gritsay et al. (2023a). More than 60,000 scientific papers in the last year alone contained the evidence of the use of machine generation Gray (2024). All of that prove that it is crucial to develop systems able to counter the misuse of artificial data and signal to the reader that the content they read is generated.
@@
Another concern is the Web, overflowing with machine-generated content, often of poor quality. Such texts contribute bias to publicly available texts on the Internet, through false facts, hallucinations and spelling errors. Given the current agenda of using texts from the Internet to train new language models, all this bias will be inadvertently added to the model. Moreover,  Villalobos et al. (2022) revealed that the human-written data will run out by 2028. That means that the training sets for language models in the future will include a large amount of generated content. Such self-consuming will result in the substantial degradation of the model’s abilities Alemohammad et al. (2023). Furthermore, the trend is evolving in such a way that human-written texts on almost any topic will be much harder to retrieve. While for texts dated even 5 years ago we are confident as the usage of generation was extremely rare, we cannot state the same for more recent texts.
@@
Therefore, detectors, capable to distinguish human-written texts from AI-generated texts, and whose detection quality can be guaranteed, are necessary for many fields. We believe one of the key factors for building reliable detectors is the high-quality artificial text collections that can be used for training and evaluation. In this paper, we would like to estimate the quality of the available generated texts from competitions and research papers. Sometimes we see that some methods from participants of the competitions reach almost perfect (up to 99.9%) metric score, meanwhile in the wild we observe a noticeable decline in performance. Such results look confusing, because the models become more and more advanced, seemingly making the detecting task more challenging, meanwhile participants of competitions still reach almost perfect scores, bringing up the question about quality of generated data in the provided datasets. Are the devised methods really good or is the data easy enough for detectors to solve the seemingly hard detection task?
Our contributions are as follows.
@@
    We systemize information about existing datasets from the research papers and competitions, dedicated to the detection of AI-generated content task.
@@ 2.1 The Task of AI-generated text detection
The task of AI-generated text detection task is generally stated as a text classification task, which means that the input is a text sequence and the output is a discrete, usually binary, class prediction. When the task is binary, the common labels are “AI” or “human”, whereas multiclass classification focusses on distinguishing several language models. The last task is usually called authorship attribution. Finally, more complex task suggests to determine the borders betweens fragments from different authors, for example between human author and some LLM author.
The approach to tackle the classification problem is to utilize linguistic, stylometric, and statistical features for classifiers Jawahar et al. (2020); Fröhling and Zubiaga (2021). Some internal metrics, such as perplexity or its modifications Hans et al. (2024) can also be helpful as a inference method. Also, perturbing texts can also provide valuable information, as described in DetectGPT Mitchell et al. (2023) method, where log-probabilities between between original and modified texts are compared. Finally, finetuning encoder-based model, like DeBERTa He et al. (2021), is considered the best solution for this task Uchendu et al. (2021); Macko et al. (2023).
2.2 Evaluating Generated Text
@@
As for evaluating the quality of the generated data itself, it has become more common to evaluate it with the help of LLMs Xu et al. (2023). This approach does not require any human reference, unlike ROUGE Lin (2004). However, the downside is the output of model-evaluator needs to be unified, not always interpretable and model-evaluator scores can be skewed. Alternative approach is suggested by  Zhu and Bhat (2020), where the text is evaluated based on several linguistic criteria, such as grammar or coherence.
2.3 Datasets with Artificial Content
There are a number of surveys about machine-generated content detection with an overview of the datasets Jawahar et al. (2020); Wu et al. (2023), however, few works focus on the quality of data in the available datasets, despite it being an important aspect of the task. Building AI-generated content detectors requires high-quality labelled data that involve substantial financial, computational, and human resources. The human evaluators should check that the dataset does not contain corrupted generations, that the texts are coherent and grammatically correct. We will describe the datasets we used in our analysis and experiments in Section 3.
2.4 Shared Tasks on AI-Generated content Detection
Shared tasks help to advance research on detecting AI-generated content, offering new variations on the tasks and data for evaluation, encouraging participants to come up with novel ideas for detectors robust to the change of language, domain, or generating model. Participants explore approaches ranging from transfer learning on complex text features to utilising and fine-tuning LLMs for these tasks (2022; 2022; 2023b; 2024; 2024; 2024; 2024; 2024). These efforts have highlighted challenges such as handling multilingual data and adapting to rapidly evolving generative models. Some participants also provide some analysis of the given data or even discuss some flaws with generated texts in it Voznyuk and Konovalov (2024).
3 Datasets
Dataset Year Language # Texts, $10^{3}$ Texts, G / H, $10^{3}$ Average Length, G / H Median Length, G / H
GPT2 2019 en 1250 1000 / 250 2941 / 2616 3245 / 2459
TweepFake 2019 en 20.7 10.4 / 10.4 104 / 118 89 / 94
HC3 2023 en, zh 85.4 26.9 / 58.5 1011 / 681 1012 / 422
GhostBuster 2023 en 21 18 / 3 3345 / 3391 3440 / 2911.5
MGTBench 2024 en 23.7 20.7 / 3 1596 / 3391 1226 / 2911.5
MAGE 2024 en 436 152.3 / 284.2 1139 / 1282 706 / 666
M4 2024 en, zh, ru, bg, ur, id 89.5 44.7 / 44.7 1588 / 3162 1454 / 1697
OutFox 2024 en 57.6 43.2 / 14.4 2686 / 2238 2311 / 1992
DAGPap22 2022 en 5.3 3.6 / 1.6 799 / 1180 680 / 1126.5
RuATD 2022 ru 129 64.5 / 64.5 237 / 221 99 / 95
AuTex 2023 en, es 65.9 33.1 / 32.8 315 / 297 386 / 351
IberAuTex 2024 es, en, ca, gl, eu, pt 98 52.5 / 45.4 1037 / 1058 981 / 1018
PAN24 2024 en 15.2 14.1 / 1.1 2641 / 3007 2731 / 2868
SemEval24 Mono 2024 en 34.2 18 / 16.2 2465 / 2358 2570 / 2083.5
SemEval24 Multi 2024 en, ar, de, it 42.3 22.1 / 20.2 2218 / 2257 2270 / 2032
MGT-1 Mono 2025 en 610.7 381.8 / 228.9 1448 / 1541 1208/ 1080
MGT-1 Multi 2025 en, zh, it, ar, de, ru, bg, ur, id 674 416.1 / 257.9 1423 / 1445 1195 / 1032
Table 1: Statistics of the texts in the datasets from the shared tasks and research papers.
@@ 3.1 Datasets From Shared Tasks
The most common tasks in shared tasks are binary classification and authorship attribution, with binary classification being the prevalent task, therefore, in this work, we focused only on it. All chosen shared tasks contain texts in English, unless stated otherwise. Here we give a brief overview of each task, as well as some quantitative statistics of the texts in Table 1, whereas a more detailed description, such as models used for generation or domains of the presented texts, can be found in Appendix A.
  1. 1.
    DAGPap 2022 Kashnitsky et al. (2022) introduced a dataset of human- and machine-written scientific excerpts collected by Elsevier.
  2. 2.
    RuATD 2022 Shamardina et al. (2022) focused on human- and machine-written documents in Russian, covering a wide range of themes.
  3. 3.
    AuTexTification 2023 Sarvazyan et al. (2023) provided texts in English and Spanish, covering five distinct domains.
  4. 4.
    IberAuTexTiification 2024 Sarvazyan et al. (2024) expanded on the previous competition with a multilingual (six Iberian languages), multi-domain, and multi-model focus.
  5. 5.
    Voight-Kampff Generative AI Authorship Verification 2024 Ayele et al. (2024), hereafter referred to as PAN 2024, tasked participants with identifying the human-authored text from two samples – one human-written and one machine-generated.
  6. 6.
    SemEval 2024 Task 8 Wang et al. (2024c) addressed domain, generator, and language shifts in generated texts. Training data included multiple languages (e.g., Chinese, Urdu, and Russian), but the test set was limited to English, Italian, German, and Arabic.
  7. 7.
    MGT Detection Task 1 (COLING 2025) Wang et al. (2025) was built on SemEval 2024 Task 8 by incorporating data generated by novel LLMs and expanding multilingual coverage of train and test sets.
@@ 3.2 Datasets from Papers
The number of collections with generated content has started to grow with an increasing number of available generators. Quite often, researchers, together with a new approach for AI content detection, publish a parallel dataset on which they have validated their method. In this paper, our aim was to pick collections with human- and machine-generated excerpts that are the most common and cited in other researchers’ publications. Similarly to previous subsection, here we give a brief overview of each chosen datasets, describe some statistics about the texts in Table 1, and add a more detailed description in Appendix A.
GPT2 Output Dataset [^1] https://github.com/openai/gpt-2-output-dataset consists of text outputs generated by GPT-2 models of different sizes across various prompts.
@@
M4 (Multilingual, Multimodal, Multitask, Massive Dataset) Wang et al. (2024b) is a large-scale dataset designed for training models that can handle multiple languages, tasks, and modalities, making it useful for developing versatile AI systems. Although it is multilingual, we sampled only English texts.
TweepFake Fagni et al. (2021) contains real tweets written by humans and synthetic tweets, generated by various AI models, from bots, imitating human users.
Outfox Koike et al. (2024) contains triplets of essay problem statements, human-written essays, and LLM-generated essays. The students who wrote the essays range from 6th to 12th grade in the USA.
Figure 1: Comparison of embedding shifts after two types of modifications for the HC3 dataset.
Figure 2: Topological Time Series for different datasets. The results for the remaining datasets selected in this paper can be found in Figure 4.
@@ 4 Approach
We decided to evaluate all datasets with common setups to see how good standard approaches perform on them. We did not have the goal to obtain the highest score, but rather to compare the performance of the same method on different datasets.
4.1 Baselines
We used several common classifiers as baselines on each dataset to obtain first-hand understanding of the dataset. The first one is mDeBERTa He et al. (2021), which is the current state-of-the-art model for multilingual machine-generated text detection Macko et al. (2023). The next baseline is DetectGPT framework Mitchell et al. (2023). We used GPT-2 Radford et al. (2019) as the base model and T5-Large Raffel et al. (2019) as perturbations generator. However, as DetectGPT requires intensive computational costs, we utilized Fast-DetectGPT Bao et al. (2024) that substitutes DetectGPT’s perturbation step with a more efficient sampling step. Finally, we used Binoculars Hans et al. (2024) as a method with improved perplexity score. The last two baselines require no fine-tuning which is crucial for detection task, as it is infeasible to train the detector on all existing artificial texts.
4.2 Topological Statistics
It was shown in Tulchinskii et al. (2023) that if we take the inner dimensionality of the manifold on the set of embeddings, we could separate human-written texts from machine-generated ones. The authors used persistence homology dimension (PHD) and showed that statistically human-generated texts have higher PHD than machine-generated texts, therefore introducing a novel detector. We calculated PHD on each set of texts. Additionally, in  Kushnareva et al. (2024) it was suggested to calculate PHD within sliding window. These intrinsic dimensions of the text within sliding window can be used as a feature for detector. The authors conclude that this metric as well can help to differentiate texts of different origin. To be able to compare datasets between each other, we came up with a score, utilizing KL-divergence, but we also needed it to be symmetrical. Let $h_{d}$, $m_{d}$ be distributions of intrinsic dimensions for two types of texts from the same dataset, of human and machine origin, then our $\text{KL}_{\text{TTS}}$ is following:
@@ 4.3 Perturbations and Shuffling
Based on the results of the text modification studies Sadasivan et al. (2024); Mitchell et al. (2023), which show how small perturbations affect machine reading comprehension systems, we decided to consider this way of possibly assessing the quality of a dataset. The key idea here is that AI models are sensitive to such adversarial changes, unlike humans. We considered two modification ideas: Adversarial Token Perturbation and Sentence Shuffling.
Adversarial Token Perturbation. In this approach we divide the text into tokens and randomly replace the token with a synonym from the WordNet collection Miller (1994) with a probability of 70%. We apply such a technique to each represented class. Using an encoder model, we obtain embeddings for each of the texts in the current dataset. Finally, we measure the average embedding shifts for the classes of human and generated texts. We obtain the embedding shifts using the cosine distance between the embeddings of the original texts and the modified ones. As a result, after modifications we obtain $\Delta_{\text{shift}}$ — the log difference of the average embedding shifts.